OpenStudy (anonymous):

A company’s new product is sold for $16.50 per unit. Product sales in the United States earn $400,000 and product sales in Europe earn $230,000. How many units does it sell altogether in the United States and Europe? a. 13,939 units c. 24,242 units b. 38,181 units d. 381,181 units

OpenStudy (anonymous):

Units sold in US= (Total Money Made in US)/(Price Per Unit)=400,000/16.5=24242 Units sold in EURO =(Total Revenue in EURO)/(Price Per Unit) = 230,000/16.5= 13939 Total Sold= 24242+13939= 38181 Units

OpenStudy (anonymous):

thank you